Hi my name is jack and I need help factoring this problem.

x^2+7x-30

thank you

Found 2 solutions by Mathtut, kvs_sandhya:
Answer by Mathtut(3670)
Answer by kvs_sandhya(2)

x^2+7x-30
Inthis equaltion,we have the co effiecient of x as +7 and the multiplied factor of co efficients of x^2 and the constant as -30
so we look for factors of 30 whose difference will give us a 7 as the result.
30 can have factors 3 and 10
3*10 =30
10-3 = 7
which satisfies our criteria.
since the product of teh figures has to be -30 one of the numbers should bear a negative sign.The difference between the numbers however is positive ,so we have 10 as a positive number and 3 as a negative
therefore
x^2-3x+10x-30 =0
x(x-3)+10(x-3)=0
(x-3)(x+10) are the factors of the equation